# AutoSAM: Towards Automatic Sampling of User Behaviors for Sequential Recommender Systems

### v3

@@ A.2.1. Optimal Hyper-parameter
Figure 7. Ablation Study of temperature.
A.2.2. Ablation Study of Temperature
The temperature $\tau$ in the Softmax function of the sampler significantly influences the distribution of sample probabilities, consequently impacting the recommendation performance of the sequential recommender system. In this part, we analyze this impact by varying $\tau$ from $\{1,3,5,7,9\}$. As shown in Figure 7, we observe that AutoSAM achieves the best performance at $\tau=5$. Deviating from this value, either by choosing too large or too small $\tau$, leads to a decline in performance. A possible explanation is that smaller $\tau$ may cause convergence to local optima, while larger $\tau$ results in nearly uniform distributions.

### v4

@@ 5.2.2 Generality Analysis
Figure 3: Performance and computation cost w.r.t. sample rates by setting relax factor from $\{-0.5,0,...,2\}$. The dashed line represents the baseline model, i.e., FullSAM.
5.2.4 Efficiency and Effectiveness of Different Relax Factors
In practice, it is necessary to balance efficiency and effectiveness when adapting to online services. Thus, we control the relax factors $b$ in Eq. 11 to solve this problem. The performance and the computation cost of the well-trained model in inference w.r.t. average sample rate by setting relax factor from $\{-0.5,0,...2\}$ are shown in Figure 3. Note that the computation cost is measured with million floating-point operations (MFLOPs). We surprisingly observe that even preserving 50% interactions can achieve the comparable performance as baseline. However, while the sample rate reaches 80%, the performance becomes to drop which probably because the sampler leads to a uniform distribution. Besides, the shortened sequences save lots computations especially with small $b$, highlighting the value in practical applications.
5.2.5 Impact of Sequence Length.
AutoSAM can model users’ non-uniform interest over interacted items, so the method should compatible with different sequence length. To verify such effectiveness, we analyze the performance improvement of AutoSAM on sequences of different lengths. As shown in Table 4. We can find that (1) AutoSAM is also effective for users with short sequences. (2) Longer sequences do maximize the effectiveness of sampling, This is probably because longer sequences not only contain richer information but also tend to include noise, offering greater potential for optimization.
5.2.6 Sampling Quality Evaluation.
Next, we evaluate the sample qualities by comparing the portion of purchasing behaviors on sampled and dropped items of different sampling strategy. Commonly, items that are purchased or added to favorites should be of more importance than clicked ones. As displayed in Figure 4, AutoSAM retains the most purchasing behavior, showing that the learned distributions are more significant to the user’s preference, which demonstrates the effectiveness of the sampler and also implies the necessity of treating behaviors non-uniformly. As it still keeps lots of clicked interactions, a possible reason is that the click items can also reflect the user’s interests to a great extent, and the factors behind that are usually very complex in the real world.
Figure 4: Portion of purchasing behaviors of sampled and discarded interactions on Tmall Dataset.
6 Conclusion
In this work, we proposed a general automatic sampling framework, named AutoSAM, to non-uniformly treat historical behaviors. Specifically, a light-weighted sampler was first leveraged to adaptively explore the distribution of raw input, so that the sequential recommender systems (SRS) could be trained with more informative and diverse samples. Considering the challenges of non-differentiable actions and the necessity to introduce multiple decision factors for sampling, we further introduced a novel reinforcement learning-based method to guide the training of the sampler in an end-to-end manner. We conducted extensive experiments on four public datasets. The experimental results showed that the AutoSAM could obtain a higher performance gain by adaptively sampling informative items. We hope this paper could inspire more works to be proposed from data perspective for SRS.
Acknowledgement This research was supported by grants from the grants of Provincial Natural Science Foundation of Anhui Province (No.2408085QF193), USTC Research Funds of the Double First-Class Initiative (No. YD2150002501), the National Natural Science Foundation of China (62337001), the Key Technologies R & D Program of Anhui Province (No. 202423k09020039) and the Fundamental Research Funds for the Central Universities (No. WK2150110032).